OpenStudy (7thgrader):

Which equations are equivalent to the given equation? Choose all answers that are correct. A. B. C. D.

OpenStudy (7thgrader):

@ganeshie8

OpenStudy (7thgrader):

0.4y + 10

OpenStudy (7thgrader):

@sammixboo @ganeshie8 @Marki @animal108 @TexasBoy123 @mickyboy0202 @Somy

OpenStudy (7thgrader):

@ElonaSushchik

OpenStudy (somy):

did you try solving it?

OpenStudy (7thgrader):

Will fan and medal, thanks for coming and mabey helping and yes i tried solving it

OpenStudy (7thgrader):

it's kinda confusing to be honest

OpenStudy (7thgrader):

i think A?

OpenStudy (7thgrader):

it's more than one btw

OpenStudy (somy):

can you show how you did? maybe there we can see where you are going wrong

OpenStudy (7thgrader):

one sec

OpenStudy (7thgrader):

what program should i use to show you?

OpenStudy (somy):

u can take a pic and 'attach file' in here

OpenStudy (7thgrader):

ok one sec

OpenStudy (7thgrader):

does A make any sense?

OpenStudy (7thgrader):

or how about B and C?

OpenStudy (7thgrader):

i dont think im doing this right :/

OpenStudy (somy):

okay look i'll do it 0.4y= 10/ 0.4 so 0.4y= 25 which means option A is wrong

OpenStudy (7thgrader):

oh

OpenStudy (somy):

option B 0.4y / 0.4 = 10 / 0.4 now we gotta make 0.4y as subject on the left side so for that we need to get rid of 0.4 of the left side and to do it we gotta move it to right side 0.4y= 10/0.4 * 0.4 as we move it to right - we multiply it by right side so final answer when you calculate it is 0.4y= 10 which fits your answer

OpenStudy (somy):

so B is right try doing the rest now

OpenStudy (7thgrader):

ok thank you dont leave i have something for you

OpenStudy (7thgrader):

@Somy

OpenStudy (7thgrader):

and i think C? also

OpenStudy (somy):

ahahahaha such a fit tho Putin - i grew up in Russia xDDD and bingo! you are right)